OpenStudy (imtiaz7):

Use factorization to find the square root of the following expression: (x^2+3x+2)(x^2+4x+3)(x^2=5x+6) @mary.rojas @xxAshxx

OpenStudy (anonymous):

Factor each trinomial separately:\[\bf x^2+3x+2=(x+1)(x+2)\]\[\bf x^2+4x+3=(x+1)(x+3)\] @Imtiaz7 What's the last trinomial in which there is an equal sign? =P

OpenStudy (imtiaz7):

the answer is (x+1)(x+2)(x+3)

OpenStudy (anonymous):

Now note assuming the '=' sign is actually '+' we get this factorisation for the last trinomial:\[\bf x^2+5x+6=(x+2)(x+3)\]Now re-write the whole product in the factored form:\[\bf (x+1)(x+2)(x+1)(x+3)(x+2)(x+3)=(x+1)^2(x+2)^2(x+3)^2\]Taking the square root we get:\[\bf \sqrt{(x+1)^2(x+2)^2(x+3)^2}=(x+1)(x+2)(x+3)\]

OpenStudy (kenljw):

(x^2+3x+2)(x^2+4x+3)(x^2=5x+6) (x + 2) (x +1) (x + 3) (x + 1) (x +2) (x +3) (x +1)^2 (x + 2)^2 (x + 3)^2

OpenStudy (anonymous):

@Imtiaz7

OpenStudy (imtiaz7):

thanks @genius12